# Rational angle bisectors on the coordinate plane and solutions of Pell’s equations

### v16

@@ Abstract
On the coordinate plane, the slopes of two straight lines and the slope of one of their angle bisectors satisfy the equation Recently, an explicit formula for non-trivial integral solutions of this equation with solutions of negative Pell’s equations was discovered by the author. In this article, for a given square-free integer and a given integer we describe every integral solution of such that are coprime by using the fundamental unit of and elements of whose absolute value of norms are the smallest prime powers. We also describe every non-trivial rational solution of the above equation as one of its applications.
@@ 1 Introduction
On the coordinate plane, the following problem can be considered.
@@
in (see [4, Proposition 1]). We say that a solution of ( ‣ 1) is trivial if
@@
Every non-trivial rational solution of ( ‣ 1) is given by
@@
For each equation we call its positive integral solution with minimum its fundamental solution.
@@
Suppose that is a square number, and let be an integral solution of If is primitive, that is, are coprime, then is strictly primitive; otherwise, there exists a common prime divisor of and dividing twice and also twice, which contradicts that are coprime and is square-free.
@@
Let be a non-trivial rational solution of ( ‣ 1).
@@
The result in Case (I) refines the previous result [4, Theorem 2]. For any rational solution of ( ‣ 1) in Case (I), each of and is a ratio of the leg lengths of some Pythagorean triangle. We find the rational solutions of ( ‣ 1) in Case (II) by using the following Theorem. This is the second main theorem of this article.
@@ 2 Rational solutions of Pell’s equations
The positive Pell’s equation has a non-trivial integral solution independently of the value of
@@
Case: Then and therefore divides
This implies and therefore
@@
Case: Then and therefore With respect to an extension of the -adic additive valuation to this implies
@@
If has a strictly primitive integral solution then defined by is a strictly primitive integral solution of ∎
@@
Unless “ and and ” then is not greater than the ideal class number of for each
@@
If and then since has no strictly primitive integral solutions for each integer by (2). Combining this fact with (1) and (2), we obtain the desired result. ∎
@@
For any ideal in is primitive if and only if the residue group is cyclic (see [1, Corollary 6.30]). Combining this fact with the Chinese remainder theorem, we obtain the desired assertion. ∎
@@
by (see [6, Theorem 2.1.4]). Thus the assertion of (1) holds.
@@
(2.5) for some satisfying (1.5), (1.6), (1.8) by (1). Put
@@ 3 Rational angle bisectors
Let be a non-trivial rational solution of ( ‣ 1), and let

### v17

@@ Abstract
On the coordinate plane, the slopes of two straight lines and the slope of one of their angle bisectors satisfy the equation Recently, an explicit formula for nontrivial integral solutions of this equation with solutions of negative Pell’s equations was discovered by the author. In this article, for a given square-free integer and a given integer we describe every integral solution of such that are coprime by using the fundamental unit of and elements of whose absolute value of norms are the smallest prime powers. We also describe every nontrivial rational solution of the above equation as one of its applications.
@@ 1 Introduction
On the coordinate plane, we consider the following problem, which can be called the rational angle bisection problem.
@@
in (see [5, Proposition 1]). We say that a solution of ( ‣ 1) is trivial if
@@
Every nontrivial rational solution of ( ‣ 1) is given by
@@
Let be positive integral solutions of We say is smaller than if
For each equation we call its minimum positive integral solution its fundamental solution.
@@
Suppose that is a square number, and let be an integral solution of If is primitive, that is, are coprime, then is strictly primitive; otherwise, there exists a common prime divisor of and dividing twice and also twice, which contradicts that are coprime and is square-free. In particular, every integral solution of is strictly primitive.
@@
Let be a nontrivial rational solution of ( ‣ 1).
@@
The result in Case (I) refines the previous result [5, Theorem 2]. For any rational solution of ( ‣ 1) in Case (I), each of and is a ratio of the leg lengths of some Pythagorean triangle. We find the rational solutions of ( ‣ 1) in Case (II) by using the following Theorem. This is the second main theorem of this article.
@@ 2 Rational solutions of Pell’s equations
The positive Pell’s equation has a nontrivial integral solution independently of the value of
@@
Proof.
See [1, Theorem 3.2.1].
By Fermat’s two squares theorem (see [8, Theorem 2.15]), can be expressed as for some if and only if has no prime factors congruent to modulo We can assume since Under these conditions, is a rational solution of
Conversely, if has a rational solution with then which implies that has no prime factors congruent to modulo by Fermat’s theorem as above, and so does ∎
@@
Case: Then This implies that divides
and therefore in Hence
@@
Case: Then and therefore in With respect to an extension of the -adic additive valuation to this implies
@@
We prove the contraposition. Suppose that has no strictly primitive integral solutions. Let be an integral solution of Then defined by is an integral solution of which implies that there exists a prime divisor of such that
Case: Then and therefore
Case: Then In this implies and therefore since Hence
Thus This implies that has no strictly primitive integral solutions. ∎
@@
Unless “ and ” then
for each
@@
If and then since has no strictly primitive integral solutions for each integer by (2). Combining this fact with (1) and (2), we obtain the desired equalities.
The desired inequality follows from the argument in the proof of (1).∎
@@
For any ideal in is primitive if and only if the residue group is cyclic (see [2, Corollary 6.30]). Combining this fact with the Chinese remainder theorem, we obtain the desired assertion. ∎
@@
by (see [7, Theorem 2.1.4]). Thus the assertion of (1) holds.
@@
(2.5) for some satisfying (1.8) and the conditions obtained by replacing with in (1.5), (1.6) because of (1). Put
@@ 3 Rational angle bisectors
Let be a nontrivial rational solution of ( ‣ 1), and let